# Can Internal Software Metrics Predict App Popularity at Launch? Yeas! and Nays!

### v1

@@ 4.2. RQ2: Can we predict an app’s number of downloads using internal software metrics?
[…]
Cumulative Distribution Function
Figure 8. CDF of DownloadsPerYear.
@@
[…]
CDF of DownloadsPerYear after Removing Outliers
Figure 9. CDF of DownloadsPerYear after removing outliers.
Next, we have three feature sets similar to the previous RQ. The Voting feature selection identified nine key features: returnQty_30, tryCatchQty_max_method, fanin_20_class, wmc_40_class, maxNestedBlocksQty_70_class, lcom_90, finalFieldsQty_95, Storage, and Photos/Media/Files.
@@
Table 9. Regression metrics (RMSE, MAE, $R^{2}$) of DownloadsPerYear regressors evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
Figure 10. CDF of absolute error for the Ridge regression model in predicting DownloadsPerYear.
[…]
CDF of absolute error for the Ridge regression model in predicting DownloadsPerYear.
To further examine model performance, we analyze the scatter plot of true versus predicted values in Figure 11. In a well-performing regression model, the data points should align closely along the diagonal, indicating accurate predictions. However, as shown in the figure, the predicted values are heavily concentrated below 20,000, despite the true values spanning a much wider range. This compression of output highlights the model’s inability to capture the full variability in DownloadsPerYear. These findings reinforce the conclusion that regression models, when relying solely on internal software metrics, lack sufficient predictive capability for this task.
Figure 11. Scatter plot of true DownloadsPerYear vs predicted DownloadsPerYear.
[…]
Scatter Plot of True DownloadsPerYear vs Predicted DownloadsPerYear
In the previous research question, internal software metrics proved ineffective for regression due to the skewed distribution of app ratings. However, for DownloadsPerYear, the limitation stems from a different issue: the target variable spans a wide range—from 0 to over 65,000—across only 381 apps, making accurate value prediction challenging. Once again, internal metrics failed to provide reliable regression performance, prompting us to reframe the problem as a classification task.
4.2.2. Classification: Predicting Popularity Based on DownloadsPerYear.
@@
The Voting feature selection approach identified five key features, publicFieldsQty_90, staticMethodsQty_95, loc_min_method, genreId_TOOLS, and Photos/Media/Files. With three feature sets: Size-only, Handpicked, and Voting, we train and evaluate the classification models. As the dataset is already balanced, resampling techniques are not required in this case. Table 10 presents the classification results, where the MLPClassifier achieved the best overall performance on the Voting feature set, with the highest MCC (0.37) and a strong AUC of 0.71. Although other models, such as Logistic Regression, attained slightly higher AUC values on specific feature sets, MLPClassifier demonstrated the most balanced performance across all three metrics within the Voting configuration.
Table 10. Classification metrics (Macro F1 Score, MCC, AUC) of DownloadsPerYear classifiers evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
The class-wise performance metrics and confusion matrix are presented in Table 11 and Figure 12, respectively. The model achieves a precision of 0.682 and a recall of 0.689 for the Unpopular category, leading to an f1-score of 0.685. Meanwhile, for the Popular category, the model attains a precision of 0.687 and a recall of 0.680, resulting in an f1-score of 0.684. These results indicate that the model maintains a balanced performance across both classes. Overall, these results again confirm the usefulness of internal metrics.
Table 11. Classification report for DownloadsPerYear-based popularity prediction using MLP classifier.
@@
[…]
Confusion Matrix and AUC-ROC Curve (DownloadsPerYear)
Figure 12. Confusion matrix and AUC-ROC curve for DownloadsPerYear-based popularity prediction using MLP classifier.
@@
5. Discussion
Our empirical investigation into the predictive power of internal software metrics for application popularity reveals both promising insights and notable limitations. In addressing RQ1—predicting application ratings via regression—we found that the models exhibited limited generalizability, reflected in low $R^{2}$ scores, and failed to capture the underlying variability in the data. However, when reframed as a classification task (i.e., distinguishing popular from unpopular apps), model performance improved, though not substantially. Still, these results suggest that internal software metrics carry meaningful predictive value at release time. A similar pattern emerged in RQ2, which examined the prediction of the DownloadsPerYear metric.
Overall, while software metrics alone may not fully explain application popularity, they offer valuable insight into its underlying dynamics. Additional non-code factors, such as developer reputation, historical performance, experience, and marketing efforts—likely play a significant role in shaping popularity outcomes (Tian et al., 2015). Incorporating such features in future work may lead to more accurate and comprehensive predictive models.
@@ 5. Discussion
Our findings suggest that different internal metrics contribute valuable signals for predicting app popularity. Practitioners should therefore view code-level metrics not only as maintainability indicators but also as potential predictors of market performance. We recommend that software development teams systematically incorporate code quality monitoring into their development processes. The integration of static analysis tools capable of tracking complexity, coupling, and code smells can aid in identifying latent issues that may impact both maintainability and user perception. Early detection of such issues allows teams to implement preventive measures, thereby improving both software quality and user satisfaction. Furthermore, practitioners are encouraged to explore the use of machine learning and predictive analytics as part of their quality assurance strategy.
@@
From a research standpoint, our findings provide meaningful insights into the role of internal software metrics in predicting application popularity. Contrary to prior studies that emphasize code size as the dominant or sole predictor of software quality and performance (El Emam et al., 2001; Herraiz et al., 2007; Gil and Lalouche, 2017; Scalabrino et al., 2016), our results demonstrate that a broader set of code metrics contribute significantly to a model’s predictive performance. This challenges the prevailing assumption that size is the only useful code metric (Gil and Lalouche, 2017; El Emam et al., 2001) and highlights the value of multidimensional code analysis. This finding should encourage the research community to develop a more comprehensive set of code metrics. Additionally, the improved performance achieved using feature selection algorithms underscores the importance of this preprocessing step in predictive modeling. These findings should motivate future work on designing and evaluating more sophisticated feature selection algorithms to further enhance prediction performance. We also observed that a model’s performance is significantly impacted by the type of algorithm used for addressing the class imbalance problem, indicating a strong need for additional research in this direction. Researchers may also focus on building genre-specific models rather than generic ones, as genre type emerged as an important feature in both of our classification models.
5.1. Threats to Validity
The external validity of our study is limited by our focus on Android apps exclusively. As a result, the findings may not be generalizable to iOS apps. Additionally, within the Android ecosystem, we concentrated solely on Java-based applications, although other programming languages, such as Kotlin, are also commonly used to develop Android apps.
@@ 5.1. Threats to Validity
6. Conclusion
This study investigates the relationship between internal software metrics and app popularity indicators, specifically user ratings and download counts. While our results indicate that internal software metrics alone are insufficient for accurately predicting continuous measures such as ratings or yearly downloads, they demonstrate substantial predictive power when app popularity is reformulated as a binary classification problem. These findings highlight the potential utility of internal code metrics in understanding and characterizing software success.

### v2

@@ 4.2 RQ2: To what extent can internal software metrics predict an app’s number of downloads?
Figure 8: CDF of DownloadsPerYear.
@@
Figure 9: CDF of DownloadsPerYear after removing outliers.
Next, we have three feature sets similar to the previous RQ. The Voting feature selection method identified nine key features: maxNestedBlocksQty_70_class, finalFieldsQty_95,Storage, wmc_40_class, maxNestedBlocksQty_70_class , Photos/Media/Files, tryCatchQty_max_method, returnQty_30, lcom_90, and fanin_20_class .
@@
Table 9: Regression metrics (RMSE, MAE, $R^{2}$) of DownloadsPerYear regressors evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
Figure 10: CDF of absolute error for the Ridge regression model in predicting DownloadsPerYear.
To further examine model performance, we analyze the scatter plot of true versus predicted values in Figure 11. In a well-performing regression model, the data points should align closely along the diagonal, indicating accurate predictions. However, as shown in the figure, the predicted values are heavily concentrated below 20,000, despite the true values spanning a much wider range. This compression of output highlights the model’s inability to capture the full variability in DownloadsPerYear. These findings reinforce the conclusion that regression models, when relying solely on internal software metrics, perform poorly.
Figure 11: Scatter plot of true DownloadsPerYear vs predicted DownloadsPerYear.
In the previous research question, internal software metrics proved ineffective for regression due to the skewed distribution of app ratings. However, in this research question for DownloadsPerYear, the limitation stems from a different issue: the target variable spans a wide range—from 0 to over 65,000—across only 381 apps, making accurate value prediction challenging. Once again, internal metrics failed to provide reliable regression performance, prompting us to reframe the problem as a classification task.
4.2.2 Classification: Predicting Popularity Based on DownloadsPerYear.
@@
The Voting method selected five features: genreId_TOOLS, loc_min_method, publicFieldsQty_90, staticMethodsQty_95, and Photos/Media/Files. With three feature sets: Size-only, Handpicked, and Voting, we train and evaluate the classification models. As the dataset is already balanced, resampling techniques are not required in this case. Table 10 presents the classification results, where the MLPClassifier achieved the best overall performance on the Voting feature set, with the highest MCC (0.37) and a strong AUC of 0.71. Although other models, such as Logistic Regression, attained slightly higher AUC values on specific feature sets, MLPClassifier demonstrated the most balanced performance across all three metrics within the Voting configuration.
Table 10: Classification metrics (Macro F1 Score, MCC, AUC) of DownloadsPerYear classifiers evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
The class-wise performance metrics and confusion matrix are presented in Table 11 and Figure 12, respectively. The model achieves a precision of 0.682 and a recall of 0.689 for the Unpopular category, leading to an F1-score of 0.685. Meanwhile, for the Popular category, the model attains a precision of 0.687 and a recall of 0.680, resulting in an F1-score of 0.684. These results indicate that the model maintains a balanced performance across both classes, unlike what we observed in RQ1 for popularity based on ratings. Overall, these results again confirm the usefulness of internal metrics.
Table 11: Classification performance for DownloadsPerYear-based popularity prediction using MLP classifier.
@@
Figure 12: Confusion matrix and AUC-ROC curve for DownloadsPerYear-based popularity prediction using MLP classifier.
@@
5 Discussion
Our empirical investigation into the predictive power of internal software metrics for application popularity reveals both promising insights and notable limitations. In addressing RQ1—predicting application ratings via regression—we found that the models exhibited limited generalizability, reflected in low $R^{2}$ scores, and failed to capture the underlying variability in the data. However, when reframed as a classification task (i.e., distinguishing popular from unpopular apps), model performance improved. These results suggest that internal software metrics carry meaningful predictive value at release time. A similar pattern emerged in RQ2, which examined the prediction of the DownloadsPerYear metric.
It is unsurprising that software metrics alone do not fully explain application popularity. Clearly, non-code factors, such as developer reputation, historical performance, the level of competition for the app, the usability of its UI, and marketing efforts, will likely play a significant role in shaping popularity outcomes (Tian et al., 2015). Incorporating these features in future work may lead to more accurate and comprehensive predictive models. Considering the potential influence of the non-code factors, it is striking that internal software metrics can still offer valuable insight into app popularity. What seems obvious from this study is that a development organization would be unwise to ignore software metrics in deciding whether a product is ready for release.
@@ 5 Discussion
Our findings suggest that different internal metrics can contribute signals for predicting app popularity. Practitioners should therefore view code-level metrics not only as maintainability indicators but also as potential predictors of market performance. We recommend that software development teams systematically incorporate code quality monitoring into their development processes. The integration of static analysis tools capable of tracking complexity, coupling, and code smells can aid in identifying latent issues that may impact both maintainability and user perception. Early detection of such issues allows teams to implement preventive measures, thereby improving both software quality and user satisfaction and potentially warding off technical debt (Cai and Kazman, 2023b). Furthermore, practitioners are encouraged to explore the use of machine learning and predictive analytics as part of their quality assurance strategy.
@@
From a research standpoint, our findings provide meaningful insights into the role of internal software metrics in predicting application popularity. Contrary to earlier studies (El Emam et al., 2001; Herraiz et al., 2007; Gil and Lalouche, 2017; Scalabrino et al., 2016) that positioned code size as the principal—or sole—predictor of software quality and performance, our findings reveal that a wider range of metrics also contribute meaningfully to predictive accuracy, especially when context is considered as the king (regression vs. classification), as was also concluded by some other studies relevant studies—class-level vs method-level evaluation (Landman et al., 2014; Chowdhury et al., 2025), and #revisions vs. change size based evaluation  (Chowdhury et al., 2022).
This finding should encourage the research community to develop a more comprehensive set of code metrics and to evaluate them based on different contexts. Additionally, the improved performance achieved using feature selection algorithms underscores the importance of this preprocessing step in predictive modeling. These findings should motivate future work on designing and evaluating more sophisticated feature selection algorithms to further enhance prediction performance. We also observed that a model’s performance is significantly impacted by the type of algorithm used for addressing the class imbalance problem, indicating a strong need for additional research in this direction. Researchers may also focus on building genre-specific models rather than generic ones, as genre type emerged as an important feature in both of our classification models. However, the small number of apps that are commonly available on both F-Droid and the Google Play Store would be a challenge for building such models.
5.1 Threats to Validity
The external validity of our study is limited by our focus on Android apps exclusively. As a result, the findings may not be generalizable to iOS apps. Additionally, within the Android ecosystem, we concentrated solely on Java-based applications, although other programming languages, such as Kotlin, are also commonly used to develop Android apps. We can not imagine any reason why the difference in platform or language would significantly affect our findings, but it still remains an external threat.
Also, our dataset is relatively small, as we specifically needed apps available on both F‑Droid and the Google Play Store. However, linking these two repositories has been the most viable approach when access to both source code and popularity indicators was required (Grano et al., 2017; Catolino, 2018; Coppola et al., 2019).
@@ 5.1 Threats to Validity
6 Conclusion
This study investigates the relationship between internal software metrics and app popularity indicators, specifically user ratings and download counts. While our results indicate that internal software metrics alone are insufficient for accurately predicting continuous measures such as ratings or yearly downloads, they demonstrate predictive power when app popularity is reformulated as a binary classification problem. These findings highlight the potential utility of internal code metrics in understanding and characterizing software success.